### v1

Abstract
Molecular structure elucidation from spectra is a foundational problem in chemistry, with profound implications for compound identification, synthesis, and drug development. Traditional methods rely heavily on expert interpretation and lack scalability. Pioneering machine learning methods have introduced retrieval-based strategies, but their reliance on finite libraries limits generalization to novel molecules. Generative models offer a promising alternative, yet most adopt autoregressive SMILES-based architectures that overlook 3D geometry and struggle to integrate diverse spectral modalities. In this work, we present DiffSpectra, a generative framework that directly infers both 2D and 3D molecular structures from multi-modal spectral data using diffusion models. DiffSpectra formulates structure elucidation as a conditional generation process. Its denoising network is parameterized by Diffusion Molecule Transformer, an SE(3)-equivariant architecture that integrates topological and geometric information. Conditioning is provided by SpecFormer, a transformer-based spectral encoder that captures intra- and inter-spectral dependencies from multi-modal spectra. Extensive experiments demonstrate that DiffSpectra achieves high accuracy in structure elucidation, recovering exact structures with 16.01% top-1 accuracy and 96.86% top-20 accuracy through sampling. The model benefits significantly from 3D geometric modeling, SpecFormer pre-training, and multi-modal conditioning. These results highlight the effectiveness of spectrum-conditioned diffusion modeling in addressing the challenge of molecular structure elucidation. To our knowledge, DiffSpectra is the first framework to unify multi-modal spectral reasoning and joint 2D/3D generative modeling for de novo molecular structure elucidation.
keywords:
Molecular Structure Elucidation, Molecular Spectra, Diffusion Models
1 Introduction
Molecular structure elucidation represents one of the fundamental challenges in chemistry, materials science, and biology [1, 2, 3], requiring the accurate determination of atomic connectivity, stereochemistry, and three-dimensional (3D) arrangements based on experimental data. Traditional approaches often require expert interpretation and substantial manual effort to propose and validate structural hypotheses. Although these methods have proven invaluable over decades of chemical research, the increasing complexity of natural products, synthetic compounds, and pharmaceutical intermediates demands more sophisticated computational approaches that can efficiently navigate the vast and complex chemical space of possible structures.
The integration of machine learning into spectroscopic analysis has opened new avenues for automating and accelerating molecular structure determination. Early applications of machine learning to molecular structure elucidation primarily adopted retrieval or search-based paradigms [1, 4, 5]. These approaches treat the task as a matching problem: given an experimental spectrum, the model searches within a predefined molecular library to retrieve candidate structures whose spectra best align with the observed one. Although efficient in practice, these methods are inherently constrained by the coverage and quality of the reference library. To mitigate this limitation, some efforts have proposed spectral prediction models, which simulate spectra from known molecular structures [6, 7, 8, 9, 10, 11, 12]. These predicted spectra are then used to expand the searchable library and improve retrieval coverage. However, such methods still rely on the existence of a finite reference library and fundamentally struggle to generalize to novel molecules that are not well represented in the library, limiting their applicability in open-ended or exploratory chemical analyses.
To overcome the reliance on molecular libraries and enable fully data-driven structure elucidation, recent research has turned to more expressive neural architectures capable of de novo structure inference directly from spectra. Initial attempts in this direction framed the task as a predictive problem, training models to infer coarse structural features from spectra, such as molecular properties [13], fingerprints [14], fragments [15, 16], chemical formulas [17], or other approximate structure profiles [18, 19, 20, 21]. Although informative, these coarse predictions do not achieve complete molecular structure elucidation.
In a shift toward complete structure elucidation, subsequent works introduced generative models that take spectral data as input and produce discrete molecular representations, such as SMILES [22] or molecular graphs [23]. These models typically follow autoregressive architectures, generating molecular structures token by token using RNNs [24, 25, 26] or Transformers [27, 16, 4, 28, 29, 30]. Despite their promise, they face two fundamental limitations. First, generative models built on discrete SMILES sequences or graphs inherently lack geometric inductive biases and cannot effectively capture 3D geometric information. This is critical because many spectral modalities, such as infrared (IR), Raman, and ultraviolet-visible (UV-Vis), are inherently dependent on the molecule’s geometry and associated energy states [31]. Second, all existing methods are limited to processing a single type of spectrum in isolation, whereas molecular structure elucidation in practice relies on the joint interpretation of multiple spectral modalities, each offering complementary structural cues [32]. Current models lack support for such multi-modal spectral reasoning, thereby constraining their accuracy and generalizability. These limitations highlight the need for a more principled and unified generative framework that can integrate both geometric reasoning and multi-modal spectral conditioning.
The inherent limitations of existing approaches motivate the development of more sophisticated generative models that can effectively bridge the gap between spectral observations and molecular structure determination. Diffusion models have recently attracted significant attention in generative modeling due to their ability to produce high-quality samples through iterative denoising processes [33, 34]. This framework has demonstrated remarkable success in modeling complex, multi-modal distributions and generating diverse outputs, and has also shown particular promise in molecular applications [35, 3]. The ability of diffusion models to generate chemically meaningful structures and handle multi-modal distributions makes them particularly well-suited for molecular structure elucidation.
In this work, we introduce DiffSpectra, a novel framework that leverages diffusion models to directly generate molecular structures from multi-modal spectra. Our approach addresses the structure elucidation problem as a conditional generation task, where spectral measurements serve as conditioning information to guide the iterative denoising process towards generating plausible molecular structures.
At the core of our framework lies DiffSpectra, a joint 2D and 3D diffusion model that simultaneously generates both molecular graphs and atomic coordinates within a unified diffusion framework. Each molecule is represented as a graph comprising node features (e.g., atom types), edge features (e.g., bond types), and 3D coordinates. This unified representation captures both the topological structure and spatial geometry. DiffSpectra formulates structure elucidation as a continuous-time conditional diffusion process, where noise is progressively added and then removed under the condition of spectral constraints. By simultaneously modeling both topological and geometric components, the model ensures that generated structures satisfy chemical bonding rules while preserving spatial consistency. This dual-space generation is essential for accurately reconstructing molecules whose spectra reflect both bonding connectivities and 3D conformations, effectively leveraging the complementary strengths of 2D topology and 3D geometry to ensure structurally consistent outputs.
To parameterize the denoising process, we introduce the Diffusion Molecule Transformer (DMT), an SE(3)-equivariant architecture designed for molecular structure generation. DMT consists of a stack of transformer-style blocks operating on three interleaved streams—node features, edge features, and coordinates—allowing for rich interaction between topological and geometric representations. Through geometry-aware multi-head attention and equivariant updates, DMT effectively models spatial relationships and maintains chemical symmetry.
To effectively condition the generation process on multi-modal spectra, we employ SpecFormer, a transformer-based spectral encoder specifically designed to process and integrate multiple types of molecular spectra simultaneously, including IR, Raman, and UV-Vis. SpecFormer partitions each spectrum into local patches, embeds them into a unified latent space, and employs multi-head self-attention to capture intra- and inter-spectrum dependencies. To enhance the quality of spectral embeddings, we pre-train SpecFormer using masked patch reconstruction and contrastive alignment with molecular structures. The resulting spectral embeddings provides powerful conditioning signals for DiffSpectra, effectively guiding generation toward candidates that are consistent with the measured spectra.
We extensively evaluate DiffSpectra to assess its effectiveness in structure elucidation from multi-modal spectra. Our results show that DiffSpectra generates chemically valid and stable molecular structures. When conditioned on spectra, our model attains a top-1 exact structure accuracy of 16.01%, and accurately reconstructs key functional groups with over 96% similarity. We further show that using a pre-trained SpecFormer significantly improves structure elucidation accuracy, and that combining multiple spectral modalities yields better performance than using any single spectrum alone. Importantly, we find that sampling multiple candidates greatly enhances the hit rate: the top-20 accuracy rises to 96.86%, indicating that correct structures are consistently ranked among the plausible outputs. This highlights the practical value of diffusion-based sampling in structure elucidation, where generating a shortlist of candidates enables reliable downstream validation. Our ablations further confirm the advantage of model-based SE(3) equivariance and demonstrate that moderate sampling temperatures strike a good balance between diversity and accuracy. These findings highlight the potential of spectrum-guided diffusion models for accurate molecular structure elucidation.
To the best of our knowledge, this is the first work to leverage diffusion models for elucidating complete 3D molecular structures directly from multi-modal spectral data.
2 Results
2.1 Overview of DiffSpectra Framework
To achieve accurate molecular structure elucidation from molecular spectra, we propose DiffSpectra, a spectrum-conditioned diffusion framework that generates molecular structures guided by spectral information. The overall paradigm is illustrated in Figure 1.
Formally, a molecule characterized by its structure and spectra is represented as $\mathcal{M}=(\mathcal{G},\mathcal{S})$. The molecular structure is modeled as a graph $\mathcal{G}=(\mathbf{H},\mathbf{A},\mathbf{X})$, where $\mathbf{H}\in\mathbb{R}^{N\times d_{1}}$ denotes node-level attributes such as atom types and charges, $\mathbf{A}\in\mathbb{R}^{N\times N\times d_{2}}$ encodes pairwise edge features including bond existence and bond types, and $\mathbf{X}\in\mathbb{R}^{N\times 3}$ corresponds to the 3D coordinates of the atoms. Here, $N$ is the number of atoms, $d_{1}$ is the node feature dimensionality, and $d_{2}$ is the edge feature dimensionality. The atomic ordering is consistent across $\mathbf{H}$ and $\mathbf{X}$ to preserve the correspondence between atom types and spatial positions.
Meanwhile, the molecular spectra are denoted as $\mathcal{S}=({\bm{s}}_{1},\ldots,{\bm{s}}_{|\mathcal{S}|})$, where $|\mathcal{S}|=3$ in our setting, covering UV-Vis, IR, and Raman spectra. Concretely, ${\bm{s}}_{1}\in\mathbb{R}^{601}$ records the UV-Vis spectrum from 1.5 to 13.5 eV with 0.02 eV intervals; ${\bm{s}}_{2}\in\mathbb{R}^{3501}$ is the IR spectrum spanning 500–4000 cm-1 at 1 cm-1 resolution; and ${\bm{s}}_{3}\in\mathbb{R}^{3501}$ is the Raman spectrum on the same range. Together, these modalities provide a comprehensive description of the molecular signatures.
Figure 1: (A) Overview of the DiffSpectra framework, illustrating the continuous-time forward diffusion process and reverse denoising process. The Diffusion Molecule Transformer (DMT) is used as the denoising network, and spectra encoded by SpecFormer serve as conditional input. (B) Architecture of the DMT, which processes node features, edge features, and atomic coordinates through three parallel streams with shared condition encoding, relational multi-head attention, and equivariant updates. (C) Architecture and pre-training strategy of SpecFormer, which encodes multi-modal spectra (UV-Vis, IR, and Raman) through a unified transformer encoder. It is pre-trained via masked patch reconstruction (MPR) and contrastive learning with 3D molecular structures.
Within DiffSpectra, a continuous-time variational diffusion model progressively perturbs the molecular graph with Gaussian noise during the forward process, then denoises it in reverse, conditioned on spectra-derived information. A dedicated backbone, the Diffusion Molecule Transformer, parameterizes the denoising process while maintaining permutation and SE(3) equivariance. DMT is guided by spectral embeddings extracted by SpecFormer, a Transformer-based encoder pre-trained with both masked reconstruction and contrastive alignment to capture spectrum–structure correlations. This design allows the framework to achieve faithful and chemically meaningful structure elucidation, supported by physically grounded diffusion and rich spectral priors. More details of the full architecture and training objectives are presented in  Section 4.
Table 1: Evaluation of basic metrics for molecular generation on the QM9S dataset. Metrics include stability, validity, uniqueness, and novelty, as well as distribution-based and 3D geometry-based metrics. The Train row reports statistics from the training set as a reference distribution. DiffSpectra is compared to unconditional molecular generative models.
Metric-2D Atom stable $\uparrow$ Mol stable $\uparrow$ V&C $\uparrow$ V&U $\uparrow$ V&U&N $\uparrow$ FCD $\downarrow$ SNN $\uparrow$ Frag $\uparrow$ Scaf $\uparrow$ Train 99.9% 98.8% 98.9% 98.9% 0.0% 0.063 0.490 0.992 0.946 CDGS [36] 99.7% 95.1% 95.1% 93.6% 89.8% 0.798 0.493 0.973 0.784 JODO [37] 99.9% 98.8% 99.0% 96.0% 89.5% 0.138 0.522 0.986 0.934 DiffSpectra 99.9% 98.2% 98.4% 96.8% 93.7% 0.088 0.531 0.992 0.943
Metric-3D Atom stable $\uparrow$ Mol stable $\uparrow$ FCD $\downarrow$ Bond $\downarrow$ Angle $\downarrow$ Dihedral $\downarrow$ Train 99.4% 95.3% 0.877 $5.44\text{e}{-4}$ $4.65\text{e}{-4}$ $1.78\text{e}{-4}$ E-NF [38] 84.7% 4.5% 4.452 0.6165 0.4203 0.0056 G-SchNet [39] 95.7% 68.1% 2.386 0.3622 0.0727 0.0042 G-SphereNet [40] 67.2% 13.4% 6.659 0.1511 0.3537 0.0129 EDM [41] 98.6% 81.7% 1.285 0.1303 0.0182 6.64e-4 MDM [42] 99.2% 89.6% 4.861 0.2735 0.0660 0.0239 GeoLDM [43] 98.9% 89.7% 1.030 0.2404 0.0100 6.59e-4 JODO [37] 99.2% 93.4% 0.885 0.1475 0.0121 6.29e-4 DiffSpectra 99.2% 92.9% 0.943 0.1792 0.0054 2.01e-4
2.2 DiffSpectra generates valid and stable molecular structures
As summarized in Table 1, our proposed DiffSpectra consistently achieves strong results in generating chemically valid and stable molecular structures under spectrum-conditioned settings. In the 2D evaluation, DiffSpectra attains 99.9% atom stability and 98.2% molecular stability, demonstrating performance comparable to the training data distribution, while outperforming or matching general molecular generative models such as CDGS [36] and JODO [37]. Moreover, DiffSpectra maintains high structural uniqueness and novelty, with V&U (valid and unique) at 96.8% and V&U&N (valid, unique, and novel) at 93.7%, significantly surpassing prior methods. In distribution-based metrics, DiffSpectra achieves the lowest FCD, indicating a close distributional match to the reference test set. It also reaches the best SNN, Frag, and Scaf scores, demonstrating a strong capacity to cover diverse chemical substructures and scaffolds while maintaining validity.
In the 3D evaluation, DiffSpectra demonstrates competitive geometric stability, with atom stability and molecular stability again comparable to the training set. Notably, DiffSpectra achieves the best angle and dihedral MMDs, suggesting accurate recovery of bond angles and torsional angles, and achieves a bond distance MMD comparable to the top-performing baselines. This highlights DiffSpectra ’s effectiveness in reconstructing not only 2D connectivity but also realistic 3D molecular conformations.
The Train row in Table 1 reports the statistics of the training dataset as a reference distribution. For some metrics that measure similarity to the test set, such as SNN, Frag, Scaf, or geometric MMDs, the training data does not necessarily achieve the best possible scores, since the test distribution may differ from the training distribution. All other baselines listed in Table 1, including CDGS [36], JODO [37], EDM [41], and others, are unconditional molecular generative models, which sample from the general molecular space without explicit external guidance. In contrast, DiffSpectra is specifically designed for molecular structure elucidation conditioned on spectra. The use of spectral signals as conditions provides external structural priors, enabling the model to generate results that better match the test set distribution. Therefore, the higher V&U, V&U&N, SNN, Frag, Scaf, and MMD metrics achieved by DiffSpectra compared to unconditional models are expected, as the conditioning naturally constrains the generation process toward structures consistent with the observed test samples.
Figure 2: Visualization of molecular structure elucidation results using DiffSpectra under different configurations. We compare single-spectrum inputs (IR, Raman, UV-Vis), multi-modal spectra, and the effect of incorporating the pre-trained SpecFormer. Ground-truth structures are shown on the left for reference.
Table 2: Structure elucidation performance of DiffSpectra on the QM9S dataset. We compare two configurations: one using a pre-trained SpecFormer as the spectral condition encoder, and one using an untrained SpecFormer. Reported metrics include exact structure recovery, graph overlap, fingerprint-based similarities, fragment-level similarity, and functional group similarity.
Model Pre-trained SpecFormer ACC@1 $\uparrow$ MCES $\downarrow$ $\operatorname{TaniSim}_{\mathrm{MG}}$uparrow $\operatorname{CosSim}_{\mathrm{MG}}$uparrow $\operatorname{TaniSim}_{\mathrm{MA}}$uparrow FraggleSim $\uparrow$ FGSim $\uparrow$ DiffSpectra 16.01% 1.3552 0.7837 0.8421 0.9227 0.9481 0.9618 14.11% 1.7795 0.7205 0.7938 0.8924 0.9383 0.9490
2.3 DiffSpectra accurately elucidates molecular structures from spectra
As shown in Table 2 and Figure 2, DiffSpectra demonstrates strong performance in the challenging task of molecular structure elucidation from spectra. First, the top-1 accuracy (ACC@1) reaches 16.01%, meaning that the model is able to recover the exact target structure among its predictions in a nontrivial fraction of cases. Beyond strict exact matches, the average MCES score is 1.3552, indicating that even when the model fails to generate an exact structure, it still preserves a substantial portion of the molecular graph connectivity.
In terms of fingerprint-based similarity metrics, DiffSpectra achieves a Tanimoto similarity of 0.7837 and a cosine similarity of 0.8421 over Morgan fingerprints, reflecting high agreement on local structural features. The Tanimoto similarity over MACCS keys is even higher at 0.9227, highlighting good coverage of predefined chemical substructures. These results suggest that DiffSpectra captures the key functional groups and local motifs characteristic of the target molecules.
Furthermore, the fragment-based Fraggle similarity reaches 0.9481, showing that large chemically meaningful fragments are well recovered. The functional group similarity (FGSim) is also high at 0.9618, confirming that the predicted molecules retain nearly all functional group types present in the ground-truth structures. Overall, these metrics collectively indicate that DiffSpectra is able to accurately elucidate molecular structures from spectra, recovering both global connectivity and local functional features with high fidelity.
2.4 Pre-trained SpecFormer facilitates more accurate structure elucidation
To investigate the effect of pre-training the spectrum encoder, we conduct an ablation study comparing DiffSpectra with and without a pre-trained SpecFormer module, as summarized in Table 2. When equipped with the pre-trained SpecFormer, DiffSpectra achieves a top-1 accuracy of 16.01%, compared to 14.11% without pre-training, indicating a clear improvement in correctly recovering the ground-truth molecular structures from spectra. Similarly, the MCES decreases from 1.7795 to 1.3552, suggesting that the molecular graphs predicted with pre-trained SpecFormer retain more of the correct connectivity of the target structures.
In addition, the fingerprint-based similarity scores also improve with pre-training. For example, the Tanimoto similarity over Morgan fingerprints rises from 0.7205 to 0.7837, and cosine similarity improves from 0.7938 to 0.8421. The Tanimoto similarity over MACCS keys also increases from 0.8924 to 0.9227. These gains indicate that the molecular structures generated with pre-trained SpecFormer more accurately capture both local substructural patterns and functional group motifs. Moreover, the Fraggle similarity improves from 0.9383 to 0.9481, and FGSim from 0.9490 to 0.9618, confirming that larger chemical fragments and functional groups are better preserved.
Overall, these results demonstrate that pre-training the SpecFormer encoder on spectral data provides beneficial inductive biases, allowing DiffSpectra to more effectively align spectral representations with molecular graph structures during conditional generation. This highlights the value of leveraging a well-trained spectrum encoder to facilitate accurate molecular structure elucidation from spectra.
Table 3: Effect of spectral modalities on structure elucidation performance. We report DiffSpectra’s results on the QM9S dataset using different types of spectra as conditional input, including IR, Raman, UV-Vis, and their combination. Metrics include exact structure recovery, graph overlap, fingerprint-based similarities, fragment-level similarity, and functional group similarity.
Model Spectral Modalities ACC@1 $\uparrow$ MCES $\downarrow$ $\operatorname{TaniSim}_{\mathrm{MG}}$uparrow $\operatorname{CosSim}_{\mathrm{MG}}$uparrow $\operatorname{TaniSim}_{\mathrm{MA}}$uparrow FraggleSim $\uparrow$ FGSim $\uparrow$ DiffSpectra All Spectra 14.11% 1.7795 0.7205 0.7938 0.8924 0.9383 0.9490 Only IR 10.97% 2.4812 0.6246 0.7188 0.8460 0.9197 0.9269 Only Raman 12.51% 2.1708 0.6778 0.7612 0.8701 0.9343 0.9315 Only UV-Vis 0.10% 8.7909 0.1556 0.2625 0.3634 0.5581 0.4567
2.5 Multi-modal spectra outperform single-modality spectra for structure elucidation
[…]
Among individual modalities, Raman spectra alone outperform IR and UV-Vis, with ACC@1 of 12.51% and reasonable similarity scores. IR spectra alone still provide useful chemical information, achieving 10.97% top-1 accuracy, while UV-Vis alone performs poorly (0.1% top-1 accuracy), reflecting its limited ability to uniquely identify molecular structures in the QM9S dataset. Overall, these results highlight that combining multiple spectra modalities provides complementary structural priors, enabling DiffSpectra to generate molecular structures with higher fidelity and consistency to ground-truth targets.
2.6 Sampling multiple candidates improves structural hit accuracy
Figure 3: Accuracy@$K$ with increasing number of sampled candidates. We evaluate the top-$K$ accuracy (ACC@$K$) as the number of generated candidates $K$ increases, comparing different variants of our model. Across all settings, Accuracy@$K$ consistently improves with larger $K$, confirming that multiple sampling significantly increases the chance of recovering the correct molecular structure.
Due to the inherently stochastic nature of diffusion models, each sampling process can produce a distinct yet plausible molecular structure. While a single sample may not always match the ground-truth structure exactly, generating multiple candidates increases the likelihood that at least one of them will align with the correct structure. This motivates the use of Top-$K$ Accuracy (ACC@$K$) as a more comprehensive evaluation metric, which is formally defined in Section C.2.
We present in Figure 3 the performance of different model variants under varying numbers of samples. Specifically, ACC@$K$ measures the proportion of test molecules for which the correct structure appears among the top $K$ generated candidates. For instance, the ACC@1 reflects the exact match rate under single-sample generation, while ACC@$K$ ($K>1$) demonstrate that allowing multiple guesses significantly boosts hit probability. Across all model variants, we observe a consistent upward trend in Accuracy@$K$ as $K$ increases. As $K$ increases, the model has more opportunities to generate a correct structure among its top-$K$ outputs. For example, the full model with all spectra and pre-trained SpecFormer improves from 16.01% at $K=1$ to 96.86% at $K=20$, showing that even a small sampling budget can significantly improve overall accuracy.
These results highlight an important property of diffusion-based molecular elucidation: even when the model may not always generate the correct structure in one shot, it maintains a strong ability to place the correct structure within a small number of plausible candidates. This aligns well with practical use cases, where a ranked list of likely structures can be provided for downstream validation or experimental testing. The effectiveness of multi-sample inference further demonstrates how spectral conditioning constrains the generative space, enabling the model to consistently place the correct structure among top candidates.
Table 4: Evaluation of different SE(3) equivariance strategies for DiffSpectra. We compare a model-based equivariant architecture to data-based equivariance approaches, with or without data augmentation, on molecular structure elucidation metrics. Here, “data aug” refers to data augmentation.
Model SE(3) Equivariance ACC@1 $\uparrow$ MCES $\downarrow$ $\operatorname{TaniSim}_{\mathrm{MG}}$uparrow $\operatorname{CosSim}_{\mathrm{MG}}$uparrow $\operatorname{TaniSim}_{\mathrm{MA}}$uparrow FraggleSim $\uparrow$ FGSim $\uparrow$ DiffSpectra model-based 14.11% 1.7795 0.7205 0.7938 0.8924 0.9383 0.9490 data-based (w/ data aug) 12.98% 1.8785 0.8882 0.7877 0.8882 0.9389 0.9456 data-based (w/o data aug) 7.47% 3.6036 0.5117 0.6294 0.7575 0.8665 0.8607
2.7 Comparing model-based and data-based SE(3) equivariance strategies
Since our model performs diffusion generation on molecular 3D structures, it is essential to ensure SE(3) equivariance so that the model’s predictions are consistent under rigid-body transformations of the input. To this end, we design and evaluate two strategies for achieving SE(3) equivariance: model-based equivariance and data-based equivariance. In the model-based design, geometric inductive biases are explicitly incorporated into the network architecture, for example through pairwise distance encoding and equivariant coordinate update mechanisms, ensuring equivariant behavior. In contrast, the data-based design does not include SE(3)-equivariant inductive biases in the model architecture. Instead, it relies on random rotation and translation data augmentation during training, encouraging the model to empirically learn equivariance from data.
As shown in Table 4, the model-based SE(3)-equivariant implementation of DiffSpectra achieves superior performance across most metrics. It attains a top-1 accuracy of 14.11% and the lowest MCES (1.7795), indicating stronger recovery of molecular graph connectivity. The data-based approach with data augmentation achieves slightly lower top-1 accuracy (12.98%) but shows competitive fingerprint-based and functional group similarities, with TanimotoMG, FraggleSim, and FGSim comparable to the model-based design. Notably, when data augmentation is removed from the data-based design, performance degrades substantially across all metrics, confirming the critical role of data augmentation in enabling the model to generalize over molecular rotations and translations. Overall, these results demonstrate that while data-based equivariance with augmentation can help achieve some level of SE(3)-aware behavior, explicitly enforcing geometric inductive biases through model-based equivariance is a more robust and effective strategy for molecular structure elucidation from spectra.
Table 5: Effect of sampling temperature on structure elucidation performance. We report DiffSpectra results with different sampling temperature coefficients $\tau$, which scale the injected noise during diffusion sampling. Moderate values of $\tau$ help balance diversity and accuracy, while extremely low or high $\tau$ degrade performance.
Model Temperature ACC@1 $\uparrow$ MCES $\downarrow$ $\operatorname{TaniSim}_{\mathrm{MG}}$uparrow $\operatorname{CosSim}_{\mathrm{MG}}$uparrow $\operatorname{TaniSim}_{\mathrm{MA}}$uparrow FraggleSim $\uparrow$ FGSim $\uparrow$ DiffSpectra $\tau$=1.2 15.45% 1.4224 0.7739 0.8348 0.9197 0.9468 0.9604 $\tau$=1.0 16.01% 1.3552 0.7837 0.8421 0.9227 0.9481 0.9618 $\tau$=0.8 16.30% 1.3643 0.7848 0.8429 0.9235 0.9491 0.9621 $\tau$=0.6 16.05% 1.4009 0.7794 0.8387 0.9220 0.9480 0.9627 $\tau$=0.4 15.86% 1.5081 0.7629 0.8261 0.9120 0.9442 0.9577 $\tau$=0.2 14.56% 1.9842 0.7005 0.7774 0.8743 0.9297 0.9323
2.8 Sampling temperature balances diversity and accuracy
Since a given molecular spectrum corresponds to a unique molecular structure, the task of molecular structure elucidation from spectra requires high accuracy in the generated results. However, diffusion models inherently introduce stochasticity during sampling due to the injection of random noise, which is essential for promoting diversity in the generated results. To better understand the trade-off between stochasticity and accuracy, we investigate how controlling the level of stochasticity affects performance. Specifically, we introduce a sampling temperature parameter $\tau$ in Eq. 6, which modulates the scale of the injected noise during the sampling process. Lower $\tau$ values reduce the influence of stochasticity, resulting in more deterministic outputs, while higher $\tau$ allows for greater exploration of the molecular space.
As shown in Table 5, we evaluate DiffSpectra under varying $\tau$ values ranging from 0.2 to 1.2. We observe that moderate temperature values ($\tau=0.8$ and $\tau=1.0$) achieve the best balance between diversity and correctness. In particular, $\tau=0.8$ obtains the highest top-1 accuracy of 16.30%, while maintaining competitive similarity metrics, including a TanimotoMG of 0.7848, CosineMG of 0.8429, and TanimotoMA of 0.9235. Extremely low temperatures (e.g., $\tau=0.2$) result in overly deterministic generations with reduced diversity and higher MCES (1.9842), while extremely high temperatures (e.g., $\tau=1.2$) also degrade performance due to excessive stochasticity.
Overall, these results suggest that selecting a moderate $\tau$ helps DiffSpectra strike an effective trade-off between generating structurally diverse candidates and maintaining high similarity to the ground-truth molecules.
3 Conclusion
In this work, we present DiffSpectra, a novel spectrum-conditioned diffusion framework for molecular structure elucidation. By integrating a continuous-time variational diffusion model with a Transformer-based molecular graph backbone, DiffSpectra is capable of generating molecular graphs with consistent 2D and 3D structures under spectral conditions. To encode spectral conditions, we introduce SpecFormer, a multi-spectrum Transformer encoder, which is further pre-trained via masked patch reconstruction and contrastive learning to align spectra with molecular structures. Through extensive experiments, we demonstrate that DiffSpectra achieves superior performance in recovering molecular structures from spectra, benefiting from the incorporation of pre-trained SpecFormer and multi-modal spectral inputs. In addition, our analysis shows that a model-based SE(3) equivariant architecture provides robust geometric consistency, while proper sampling temperature helps balance stochasticity and accuracy in generation.
Our work paves the way for several promising directions for future research. Firstly, expanding the diversity and quantity of molecular spectra data could further boost the performance of spectral-conditioned generation. Secondly, extending the framework to handle even more complex spectra modalities, such as NMR or mass spectra, could generalize the method toward broader applications in analytical chemistry. Lastly, adapting DiffSpectra to larger biomolecular or materials systems may open opportunities for molecular identification in drug discovery and materials science, especially when combined with high-throughput experimental pipelines. We believe DiffSpectra represents a significant step toward trustworthy and accurate molecular structure elucidation conditioned on spectral data.
4 Methods
4.1 Diffusion Model for Molecular Structure Elucidation
We adopt a continuous-time diffusion probabilistic model to generate molecular graphs in a joint space that integrates 2D topology and 3D geometry. Each molecule is represented as a graph $\mathcal{G}=(\mathbf{H},\mathbf{A},\mathbf{X})$, where $\mathbf{H}\in\mathbb{R}^{N\times d_{1}}$ denotes node-level attributes such as atom types and charges, $\mathbf{A}\in\mathbb{R}^{N\times N\times d_{2}}$ encodes pairwise edge features such as bond existence and bond types, and $\mathbf{X}\in\mathbb{R}^{N\times 3}$ corresponds to the 3D coordinates of the atoms. Here, $N$ represents the number of atoms in the molecule, $d_{1}$ denotes the dimensionality of the node features, and $d_{2}$ denotes the dimensionality of the edge features.
We adopt the mathematical formulation of the Variational Diffusion Model (VDM) [44, 45, 46] to define our diffusion and sampling processes, which follow a continuous-time diffusion probabilistic framework, as illustrated in Figure 1(A). To streamline the presentation, we consider a generic graph component, namely the node features $\mathbf{H}$, the edge features $\mathbf{A}$, or the atomic coordinates $\mathbf{X}$, which we uniformly denote by a vector-valued variable $\mathbf{G}\in\mathbb{R}^{d}$. The subsequent formulas can then be applied independently to each component in a consistent manner.
4.1.1 Forward Diffusion Process
In the forward diffusion process, Gaussian noise $\mathbf{G}_{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is gradually added to the data over continuous time from $t=0$ to $t=1$. The noised sample $\mathbf{G}_{t}$ is obtained by:
$\mathbf{G}_{t}=\alpha_{t}\mathbf{G}_{0}+\sigma_{t}\mathbf{G}_{\epsilon},$ (1)
where $\alpha_{t}$ and $\sigma_{t}$ are signal and noise scaling functions, typically determined by cosine or linear schedules. The signal-to-noise ratio (SNR) at time $t$ is defined as:
$\mathrm{SNR}(t)=\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}}.$ (2)
The SNR is strictly decreasing from $t=0$ to $t=1$, ensuring that the data is gradually corrupted into pure noise as $t\to 1$. Given two times $0\leq s<t\leq 1$, the conditional distribution of a noised graph at time $t$ given a less-noised graph at time $s$, denoted $q(\mathbf{G}_{t}\mid\mathbf{G}_{s})$, is also Gaussian:
$q(\mathbf{G}_{t}\mid\mathbf{G}_{s})=\mathcal{N}(\alpha_{t|s}\mathbf{G}_{s},\sigma_{t|s}^{2}\mathbf{I}),$ (3)
where the intermediate scaling factor and conditional variance are defined as:
$\alpha_{t|s}=\frac{\alpha_{t}}{\alpha_{s}},\quad\sigma_{t|s}^{2}=\sigma_{t}^{2}-\alpha_{t|s}^{2}\sigma_{s}^{2}.$ (4)
This formulation allows the forward process to be implemented using analytically tractable Gaussian transitions, which also facilitates efficient computation of training objectives and reverse-time sampling.
4.1.2 Reverse Denoising Process
The reverse-time denoising process starts from a standard Gaussian sample $\mathbf{G}_{1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and proceeds backward from $t=1$ to $t=0$. Let $0\leq s<t\leq 1$ denote two consecutive time steps. At each step, a data prediction model $d_{\theta}$ takes as input the noised sample $\mathbf{G}_{t}$, a self-conditioning [47] estimate $\tilde{\mathbf{G}}_{0}$, and the log signal-to-noise ratio $\log\mathrm{SNR}(t)$. The model predicts $\hat{\mathbf{G}}_{0}=d_{\theta}(\mathbf{G}_{t},\tilde{\mathbf{G}}_{0},\log\mathrm{SNR}(t))$, which is then used to compute the denoised sample $\mathbf{G}_{s}$ by:
$\bar{\mathbf{G}}_{s}=\frac{\alpha_{t|s}\sigma_{s}^{2}}{\sigma_{t}^{2}}\mathbf{G}_{t}+\frac{\alpha_{s}\sigma_{t|s}^{2}}{\sigma_{t}^{2}}\hat{\mathbf{G}}_{0},$ (5)
$\mathbf{G}_{s}=\bar{\mathbf{G}}_{s}+\tau\cdot\frac{\sigma_{s}\sigma_{t|s}}{\sigma_{t}}\mathbf{G}_{\epsilon},\quad\mathbf{G}_{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (6)
where $\tau$ is a sampling temperature parameter, modulating the amount of stochasticity in the sampling process. Lower values of $\tau<1$ reduce the influence of noise, leading to more deterministic outputs. The self-conditioning mechanism, where the model reuses its previous prediction $\tilde{\mathbf{G}}_{0}$ as an additional input in the next step, enhances training stability and generation quality [47].
4.1.3 Training Objective
To recover the original sample, we train a data prediction model $d_{\theta}$ that takes as input the noised graph $\mathbf{G}_{t}$, a self-conditioning estimate $\tilde{\mathbf{G}}_{0}$, and the log signal-to-noise ratio $\log\mathrm{SNR}(t)$. The model predicts the clean sample $\hat{\mathbf{G}}_{0}$, where $\hat{\mathbf{G}}_{0}$ refers to any component of $\hat{\mathcal{G}}_{0}=(\hat{\mathbf{H}},\hat{\mathbf{A}},\hat{\mathbf{X}})$. Therefore, the model is trained to minimize a weighted mean squared error:
$\mathcal{L}=\mathbb{E}_{t,\mathbf{G}_{0}}\left[\sqrt{\frac{\alpha_{t}}{\sigma_{t}}}\left(\lambda_{1}\|\hat{\mathbf{A}}-\mathbf{A}_{0}\|_{2}^{2}+\lambda_{2}\|\hat{\mathbf{X}}-\hat{\mathbf{X}}_{0}\|_{2}^{2}+\lambda_{3}\|\hat{\mathbf{H}}-\mathbf{H}_{0}\|_{2}^{2}\right)\right],$ (7)
where $\hat{\mathbf{X}}_{0}$ is obtained by aligning $\mathbf{X}_{0}$ to $\mathbf{X}_{t}$ using the Kabsch algorithm [48] to preserve SE(3)-equivariance. The loss coefficients $\lambda_{i}$ balance the relative importance of each component.
4.2 Diffusion Molecule Transformer (DMT)
To parameterize the data prediction module $d_{\theta}$, we employ a specialized architecture named Diffusion Molecule Transformer (DMT) [37, 49] as its backbone, to jointly model the noisy molecular graph’s node features, edge features, and 3D coordinates in our diffusion framework. The architecture is motivated by the need to recover the correlations among these three components, which are independently corrupted by noise during the forward diffusion process. DMT is specifically tailored for molecular generation tasks and respects both permutation equivariance and SE(3)-equivariance.
At each time $t\in[0,1]$ during the reverse process, DMT takes the noisy molecular graph $\mathcal{G}_{t}=(\mathbf{H}_{t},\mathbf{A}_{t},\mathbf{X}_{t})$ as input, along with a self-conditioned estimate $\tilde{\mathcal{G}}_{0}$, and predicts a clean graph $\hat{\mathcal{G}}_{0}=(\hat{\mathbf{H}},\hat{\mathbf{A}},\hat{\mathbf{X}})$. The model performs denoising inference conditioned on both the current time step and the molecular spectra. The overall model architecture is illustrated in Figure 1(B). For the sake of clarity, the self-conditioning mechanism is omitted in the figure.
A time embedding, obtained via sinusoidal encoding of $\log(\alpha_{t}^{2}/\sigma_{t}^{2})$, is passed through learnable projections. Simultaneously, a spectra embedding is generated by the SpecFormer, which will be detailed in Section 4.3. These two embeddings are then concatenated to form the conditioning vector $\mathbf{C}$, which is used throughout the network.
Initial embeddings. The DMT module is composed of stacked Transformer-style equivariant blocks. In the first layer, the node features $\mathbf{H}_{t}$ and the edge features $\mathbf{A}_{t}$ are first concatenated with their corresponding self-conditioning estimates $\tilde{\mathbf{H}}_{0}$ and $\tilde{\mathbf{A}}_{0}$, respectively. For edge features, distance features $\mathbf{D}_{0}$ computed from the self-conditioned coordinates $\tilde{\mathbf{X}}_{0}$ are also concatenated. These concatenated features are then passed through linear projections to obtain the initial embeddings:
$\mathbf{H}^{(1)}=\mathrm{Linear}_{\mathrm{node}}\bigl{(}\left[\mathbf{H}_{t},\tilde{\mathbf{H}}_{0}\right]\bigr{)}\in\mathbb{R}^{N\times d_{h}},$ (8)
$\mathbf{A}^{(1)}=\mathrm{Linear}_{\mathrm{edge}}\bigl{(}\left[\mathbf{A}_{t},\tilde{\mathbf{A}}_{0},\mathbf{D}_{0}\right]\bigr{)}\in\mathbb{R}^{N\times N\times d_{a}},$ (9)
[…]
Three-stream update. Each DMT block consists of three interacting streams that are responsible for updating node features, edge features, and coordinates, respectively. These streams exchange information throughout the block. In the node stream, information is propagated across the molecular graph via a relational multi-head attention (MHA) mechanism [50, 51], enabling flexible message passing over fully connected molecular graphs. In particular, pairwise geometric distances computed from coordinates are transformed into radial basis encodings $\rho_{ij}^{(l)}$, and these are combined with edge features and geometric distances such as:
$\bar{\mathbf{A}}_{ij}^{(l)}=\left[\mathbf{A}_{ij}^{(l)};\,\|\mathbf{X}^{(l)}_{i}-\mathbf{X}^{(l)}_{j}\|_{2};\,\rho_{ij}^{(l)}\right],$ (10)
forming a geometry-aware relational representation. The queries, keys, and values for MHA are obtained by applying learnable linear projections to node features:
$\mathbf{Q}=\mathbf{H}^{(l)}\mathbf{W}^{Q},\quad\mathbf{K}=\mathbf{H}^{(l)}\mathbf{W}^{K},\quad\mathbf{V}=\mathbf{H}^{(l)}\mathbf{W}^{V},$ (11)
where $\mathbf{W}^{Q}$, $\mathbf{W}^{K}$, and $\mathbf{W}^{V}$ are learned projection matrices. The attention weights are then defined as:
$a_{ij}=\frac{\tanh\bigl{(}\phi_{0}(\bar{\mathbf{A}}_{ij}^{(l)})\bigr{)}\,\mathbf{Q}_{i}\,\mathbf{K}_{j}^{\top}}{\sqrt{d_{k}}},\quad a=\mathrm{softmax}(a),$ (12)
and the node-level aggregation proceeds as:
$\mathrm{Attn}(\mathbf{H}^{(l)},\bar{\mathbf{A}}^{(l)})_{i}=\sum_{j=1}^{N}a_{ij}\,\tanh\bigl{(}\phi_{1}(\bar{\mathbf{A}}_{ij}^{(l)})\bigr{)}\,\mathbf{V}_{j},$ (13)
with $\phi_{0}$ and $\phi_{1}$ as learnable projections. We extend the above attention mechanism to a multi-head framework, yielding the aggregated representation $\mathbf{M}^{(l)}\in\mathbb{R}^{N\times d_{m}}$, which serves as the intermediate node embeddings after relational multi-head attention.
To incorporate time-dependent and spectra-dependent conditioning, we apply adaptive layer normalization (AdaLN) to each stream, conditioned on a learned conditioning embedding $\mathbf{C}$: $\mathrm{AdaLN}({\bm{h}},\mathbf{C})=\left(1+\mathrm{FFN}_{\mathrm{scale}}(\mathbf{C})\right)\cdot\mathrm{LN}({\bm{h}})+\mathrm{FFN}_{\mathrm{bias}}(\mathbf{C})$, and use an adaptive scaling function: $\mathrm{Scale}({\bm{h}},\mathbf{C})=\mathrm{FFN}_{\mathrm{scale}}^{\prime}(\mathbf{C})\cdot{\bm{h}}$ to modulate feature amplitudes. The node update is performed by:
$\mathbf{H}^{(l+1)^{\prime}}=\mathrm{Scale}\bigl{(}\mathbf{M}^{(l)},\mathbf{C}\bigr{)}+\mathbf{H}^{(l)},$ (14)
$\mathbf{H}^{(l+1)}=\mathrm{Scale}\left(\mathrm{FFN}\bigl{(}\mathrm{AdaLN}(\mathbf{H}^{(l+1)^{\prime}},\mathbf{C})\bigr{)},\mathbf{C}\right)+\mathbf{H}^{(l+1)^{\prime}}.$ (15)
Similarly, edge features are updated by first fusing node messages:
$\hat{\mathbf{A}}_{ij}^{(l)}=\left(\mathbf{M}^{(l)}_{i}+\mathbf{M}^{(l)}_{j}\right)\mathbf{W}_{1},$ (16)
with subsequent normalization and scaling:
$\mathbf{A}^{(l+1)^{\prime}}_{ij}=\mathrm{Scale}\bigl{(}\hat{\mathbf{A}}_{ij}^{(l)},\mathbf{C}\bigr{)}+\mathbf{A}^{(l)}_{ij},$ (17)
$\mathbf{A}^{(l+1)}_{ij}=\mathrm{Scale}\left(\mathrm{FFN}\bigl{(}\mathrm{AdaLN}(\mathbf{A}^{(l+1)^{\prime}}_{ij},\mathbf{C})\bigr{)},\mathbf{C}\right)+\mathbf{A}^{(l+1)^{\prime}}_{ij}.$ (18)
The parameter $\mathbf{W}_{1}$ is a learnable weight matrix that couples node and edge information.
For the coordinate stream, equivariant updates are achieved using directional vector fields, combining signals from node and edge streams:
$e_{ij}^{(l+1)}=\mathrm{AdaLN}\left(\mathbf{W}_{2}\left[\mathbf{H}_{i}^{(l+1)},\mathbf{H}_{j}^{(l+1)},\mathbf{A}_{ij}^{(l+1)},\|\mathbf{X}^{(l)}_{i}-\mathbf{X}^{(l)}_{j}\|_{2}\right],\mathbf{C}\right),$ (19)
$\mathbf{X}_{i}^{(l+1)}=\mathbf{X}^{(l)}_{i}+\sum_{j\neq i}\gamma^{(l)}\,\frac{\mathbf{X}^{(l)}_{i}-\mathbf{X}^{(l)}_{j}}{\|\mathbf{X}^{(l)}_{i}-\mathbf{X}^{(l)}_{j}\|_{2}}\,\tanh\bigl{(}\mathrm{FFN}(e_{ij}^{(l+1)})\bigr{)},$ (20)
where $\mathbf{W}_{2}$ is a learned projection and $\gamma^{(l)}$ is a trainable scalar that stabilizes the directional updates. Finally, coordinates are shifted to have zero center-of-mass by mean centering, ensuring translation invariance.
Overall, the DMT stacks $L$ such equivariant blocks, which collectively refine the graph features and spatial coordinates. Final output heads predict discrete atom and bond types as well as coordinates aligned to the canonical centered frame. Its design guarantees permutation equivariance and SE(3) equivariance, enabling robust molecule generation across both topological and geometric levels.
4.3 Spectra Transformer (SpecFormer) for Spectra Encoding
To enable DMT to incorporate molecular spectral information as a conditional input for molecular structure elucidation, we propose the Spectra Transformer (SpecFormer) [31] to encode molecular spectra. We further pre-train SpecFormer using both molecular structure data and spectra data.
4.3.1 Architecture of SpecFormer
For each type of spectrum, we first segment the data independently into patches and encode these patches. The patch embeddings from all spectra are then concatenated and collectively processed by a Transformer-based encoder.
Patching. Rather than encoding each frequency point individually, we divide each spectrum into several patches. This design choice offers two key benefits: (i) It allows the model to capture local semantic features—such as absorption peaks—more effectively by grouping adjacent frequency points; and (ii) it lowers the computational cost for the following Transformer layers. Formally, a spectrum ${\bm{s}}_{i}\in\mathbb{R}^{L_{i}}$ (where $i=1,\dots,|\mathcal{S}|$) is split into patches of length $P_{i}$ with stride $D_{i}$. If $0<D_{i}<P_{i}$, patches overlap by $P_{i}-D_{i}$ points; if $D_{i}=P_{i}$, patches are non-overlapping. The patching process yields a sequence ${\bm{p}}_{i}\in\mathbb{R}^{N_{i}\times P_{i}}$, where $N_{i}=\left\lfloor\frac{L_{i}-P_{i}}{D_{i}}\right\rfloor+1$ is the number of patches produced from ${\bm{s}}_{i}$.
Patch and position encoding. Each patch sequence from the $i$-th spectrum is projected into a latent space of dimension $d$ via a learnable linear transformation $\mathbf{W}_{i}\in\mathbb{R}^{P_{i}\times d}$. To preserve the sequential order, a learnable position encoding $\mathbf{W}_{i}^{\text{pos}}\in\mathbb{R}^{N_{i}\times d}$ is added: ${\bm{p}}_{i}^{\prime}={\bm{p}}_{i}\mathbf{W}_{i}+\mathbf{W}_{i}^{\text{pos}}$, yielding the encoded representation ${\bm{p}}_{i}^{\prime}\in\mathbb{R}^{N_{i}\times d}$ for each spectrum. These representations serve as input to the SpecFormer encoder.
SpecFormer: multi-spectrum Transformer encoder. Existing encoders such as CNN-AM [52] utilize one-dimensional convolutions and are typically designed to process a single type of spectrum. In contrast, our model simultaneously considers multiple molecular spectra (e.g., UV-Vis, IR, Raman) as input. This multi-spectrum strategy is motivated by the presence of both intra-spectrum dependencies—relationships among peaks within the same spectrum—and inter-spectrum dependencies—correlations between peaks across different spectral modalities. These dependencies have been well documented, for example, in studies of vibronic coupling [53].
To effectively model these dependencies, we concatenate the encoded patch sequences from all spectra: $\hat{{\bm{p}}}={\bm{p}}_{1}^{\prime}\|\cdots\|{\bm{p}}_{|\mathcal{S}|}^{\prime}\in\mathbb{R}^{(\sum_{i=1}^{|\mathcal{S}|}N_{i})\times d}$. This combined sequence is then passed through a Transformer encoder (see Figure 1). In each attention head $h=1,\ldots,H$, queries, keys, and values are computed as $\mathbf{Q}_{h}=\hat{{\bm{p}}}\mathbf{W}_{h}^{Q}$, $\mathbf{K}_{h}=\hat{{\bm{p}}}\mathbf{W}_{h}^{K}$, and $\mathbf{V}_{h}=\hat{{\bm{p}}}\mathbf{W}_{h}^{V}$ respectively, with $\mathbf{W}_{h}^{Q},\mathbf{W}_{h}^{K}\in\mathbb{R}^{d\times d_{k}}$ and $\mathbf{W}_{h}^{V}\in\mathbb{R}^{d\times\frac{d}{H}}$. The attention output for each head is:
$\mathbf{O}_{h}=\text{Attention}(\mathbf{Q}_{h},\mathbf{K}_{h},\mathbf{V}_{h})=\text{Softmax}\left(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{h}.$ (21)
Batch normalization, feed-forward networks, and residual connections are integrated as illustrated in Figure 1. The outputs from all attention heads are combined to form ${\bm{z}}\in\mathbb{R}^{(\sum_{i=1}^{|\mathcal{S}|}N_{i})\times d}$. Finally, a flattening layer and a projection head are applied to produce the molecular spectra representation ${\bm{z}}_{s}\in\mathbb{R}^{d}$.
4.3.2 Pre-training of SpecFormer
To enable more effective encoding of molecular spectra, we introduce a masked reconstruction pre-training objective. In addition, to overcome the scarcity of spectral pre-training data by leveraging large-scale molecular structure pre-training, we incorporate a contrastive learning objective to align spectral and structural representations. The complete pre-training framework is illustrated in Figure 1(C).
Masked patches reconstruction pre-training for spectra. To ensure that the spectrum encoder can effectively extract and represent information from molecular spectra, we adopt a masked patches reconstruction (MPR) pre-training strategy. Inspired by the effectiveness of masked reconstruction across multiple domains [54, 55, 56, 57, 58, 59], MPR guides the learning of SpecFormer by encouraging it to reconstruct masked portions of spectral data.
After segmenting the spectra into patches, we randomly mask a fraction of these patches—according to a predefined ratio $\alpha$—by replacing them with zero vectors. The masked patch sequences then undergo patch and position encoding, which conceals their semantic content (such as absorption intensities at certain wavelengths) but retains their positional information, aiding the reconstruction task.
Once processed by SpecFormer, the encoded representations corresponding to the masked patches are passed through a reconstruction head specific to each spectrum. The original values of the masked patches are then predicted, with the mean squared error (MSE) between the reconstructed and true patch values serving as the training objective:
$\mathcal{L}_{\mathrm{MPR}}=\sum_{i=1}^{|\mathcal{S}|}\mathbb{E}_{p_{i,j}\in\widetilde{\mathcal{P}}_{i}}\|\hat{{\bm{p}}}_{i,j}-{\bm{p}}_{i,j}\|_{2}^{2},$ (22)
where $\widetilde{\mathcal{P}}_{i}$ is the set of masked patches for the $i$-th type of spectrum, and $\hat{{\bm{p}}}_{i,j}$ denotes the reconstruction for the masked patch ${\bm{p}}_{i,j}$.
Contrastive learning between 3D structures and spectra. To align spectral and 3D molecular representations, we introduce a contrastive learning objective in addition to the MPR. The 3D embeddings are learned under the guidance of a denoising objective [60, 31]. Here, the 3D embedding ${\bm{z}}_{x}\in\mathbb{R}^{d}$ and spectra embedding ${\bm{z}}_{s}\in\mathbb{R}^{d}$ of the same molecule are treated as positive pairs, while all other pairings are considered negative. The contrastive objective is designed to maximize similarity between positive pairs while minimizing similarity with negatives, using the InfoNCE loss [61]:
$\mathcal{L}_{\text{Contrast}}=-\frac{1}{2}\mathbb{E}_{p({\bm{z}}_{x},{\bm{z}}_{s})}[\log\frac{\exp(f_{x}({\bm{z}}_{x},{\bm{z}}_{s}))}{\exp(f_{x}({\bm{z}}_{x},{\bm{z}}_{s}))+\sum_{j}\exp(f_{x}({\bm{z}}_{x}^{j},{\bm{z}}_{s}))}$ (23)
$+\log\frac{\exp(f_{s}({\bm{z}}_{s},{\bm{z}}_{x}))}{\exp(f_{s}({\bm{z}}_{s},{\bm{z}}_{x}))+\sum_{j}\exp(f_{s}({\bm{z}}_{s}^{j},{\bm{z}}_{x}))}],$
where ${\bm{z}}_{x}^{j}$ and ${\bm{z}}_{s}^{j}$ denote negative samples, and $f_{x}({\bm{z}}_{x},{\bm{z}}_{s})$ and $f_{s}({\bm{z}}_{s},{\bm{z}}_{x})$ are scoring functions, implemented here as the inner product: $f_{x}({\bm{z}}_{x},{\bm{z}}_{s})=f_{s}({\bm{z}}_{s},{\bm{z}}_{x})=\langle{\bm{z}}_{x},{\bm{z}}_{s}\rangle$.
Two-stage pre-training pipeline. Although spectral datasets are scarce, there exists a wealth of large-scale unlabeled molecular structure datasets. By leveraging our proposed contrastive alignment framework between spectra and structures, we can transfer knowledge from large-scale pre-training on molecular structures to enhance the learning of SpecFormer. To fully exploit both spectral and structural information, we adopt a two-stage training protocol: the first stage performs pre-training on a large dataset [62] containing only 3D structures using the denoising objective, while the second stage utilizes a dataset with available spectra to jointly optimize the overall objective:
$\mathcal{L}=\beta_{\text{Denoising}}\mathcal{L}_{\text{Denoising}}+\beta_{\text{MPR}}\mathcal{L}_{\text{MPR}}+\beta_{\text{Contrast}}\mathcal{L}_{\text{Contrast}},$ (24)
where $\beta_{\text{Denoising}}$, $\beta_{\text{MPR}}$, and $\beta_{\text{Contrast}}$ are weights for each component. In the second stage, SpecFormer is pre-trained exclusively on the QM9S training set to avoid data leakage.
Appendix
Appendix A Derivation of the Reverse Sampling Process
This appendix derives the reverse sampling formula for our continuous-time diffusion model, as applied to molecular graph generation. Our objective is to obtain an expression for sampling a less-noisy graph state $\mathcal{G}_{s}$ from a more-noisy state $\mathcal{G}_{t}=(\mathbf{H}_{t},\mathbf{A}_{t},\mathbf{X}_{t})$, given a model prediction of the original clean graph $\hat{\mathcal{G}}_{0}=(\hat{\mathbf{H}}_{0},\hat{\mathbf{A}}_{0},\hat{\mathbf{X}}_{0})$. We assume a continuous time interval $[0,1]$, with $0\leq s<t\leq 1$.
To streamline the derivation, we consider a single graph component—either node features $\mathbf{H}$, edge features $\mathbf{A}$, or atomic coordinates $\mathbf{X}$—denoted generically as a vector-valued variable $\mathbf{G}\in\mathbb{R}^{d}$. The full derivation extends naturally by applying the result independently to each component.
A.1 Forward Diffusion Process
Following the formulation of VDM [44, 45, 46], the forward diffusion process is defined as:
$q(\mathbf{G}_{t}\mid\mathbf{G}_{0})=\mathcal{N}(\alpha_{t}\mathbf{G}_{0},\sigma_{t}^{2}\mathbf{I}),$ (25)
where the graph $\mathbf{G}_{t}$ at time $t$ is a noisy version of the clean graph $\mathbf{G}_{0}$. Equivalently, it can be reparameterized as:
$\mathbf{G}_{t}=\alpha_{t}\mathbf{G}_{0}+\sigma_{t}\mathbf{G}_{\epsilon},\quad\mathbf{G}_{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (26)
Because $\alpha_{t}$ decreases monotonically while $\sigma_{t}$ increases, the information from $\mathbf{G}_{0}$ is gradually destroyed as $t$ increase. Assuming the process defined by Eq. 25 is Markovian, its transition distribution between two intermediate steps $s$ and $t$ with $0\leq s<t$ is:
$q(\mathbf{G}_{t}\mid\mathbf{G}_{s})=\mathcal{N}(\alpha_{t|s}\mathbf{G}_{s},\sigma_{t|s}^{2}\mathbf{I}),$ (27)
where $\alpha_{t|s}=\frac{\alpha_{t}}{\alpha_{s}}$ and $\sigma_{t|s}^{2}=\sigma_{t}^{2}-\alpha_{t|s}^{2}\sigma_{s}^{2}$. A convenient property of this framework is that the time grid can be defined arbitrarily and does not depend on the particular spacing of $s$ and $t$. We set $T=1$ to denote the final diffusion step, where $q(\mathbf{G}_{T}\mid\mathbf{G}_{0})\approx\mathcal{N}(\mathbf{0},\mathbf{I})$ approximates a standard normal distribution. Unless otherwise specified, time steps are assumed to lie in the unit interval $[0,1]$. This formulation describes the distribution of a more-noised state $\mathbf{G}_{t}$ conditioned on a less-noised state $\mathbf{G}_{s}$, and serves as a key component of the variational framework.
A.2 Reverse Denoising Process Posterior $q(\mathbf{G}_{s}\mid\mathbf{G}_{t},\mathbf{G}_{0})$
We now derive the posterior distribution $q(\mathbf{G}_{s}\mid\mathbf{G}_{t},\mathbf{G}_{0})$, which describes the distribution of an intermediate state $\mathbf{G}_{s}$ conditioned on both the noisy future $\mathbf{G}_{t}$ and the original clean graph $\mathbf{G}_{0}$. This distribution underlies the training of the reverse process.
From standard Bayesian inference for Gaussians, suppose prior ${\bm{z}}\sim\mathcal{N}(\bm{\mu}_{p},\mathbf{\Sigma}_{p})$ and likelihood ${\bm{y}}\mid{\bm{z}}\sim\mathcal{N}(\mathbf{A}{\bm{z}},\mathbf{\Sigma}_{l})$. Then the posterior $p({\bm{z}}\mid{\bm{y}})$ is Gaussian with:
$\mathbf{\Sigma}_{\text{post}}=\left(\mathbf{\Sigma}_{p}^{-1}+\mathbf{A}^{\top}\mathbf{\Sigma}_{l}^{-1}\mathbf{A}\right)^{-1},$ (28)
$\bm{\mu}_{\text{post}}=\mathbf{\Sigma}_{\text{post}}\left(\mathbf{\Sigma}_{p}^{-1}\bm{\mu}_{p}+\mathbf{A}^{\top}\mathbf{\Sigma}_{l}^{-1}{\bm{y}}\right).$ (29)
In our case:
${\bm{z}}=\mathbf{G}_{s},\quad{\bm{y}}$ $=\mathbf{G}_{t},\quad\bm{\mu}_{p}=\alpha_{s}\mathbf{G}_{0},\quad\mathbf{\Sigma}_{p}=\mathbf{\sigma}_{s}^{2}\mathbf{I},$ (30)
$\mathbf{A}$ $=\alpha_{t|s}\mathbf{I},\quad\mathbf{\Sigma}_{l}=\sigma_{t|s}^{2}\mathbf{I}.$
Applying this, we obtain:
$q(\mathbf{G}_{s}\mid\mathbf{G}_{t},\mathbf{G}_{0})=\mathcal{N}(\bm{\mu}_{Q},\sigma_{Q}^{2}\mathbf{I}),$ (31)
with:
$\bm{\mu}_{Q}=\sigma_{Q}^{2}\left(\frac{1}{\sigma_{s}^{2}}\alpha_{s}\mathbf{G}_{0}+\frac{\alpha_{t|s}}{\sigma_{t|s}^{2}}\mathbf{G}_{t}\right)=\frac{\alpha_{t|s}\sigma_{s}^{2}}{\sigma_{t}^{2}}\mathbf{G}_{t}+\frac{\alpha_{s}\sigma_{t|s}^{2}}{\sigma_{t}^{2}}\mathbf{G}_{0},$ (32)
$\sigma_{Q}^{2}=\left(\frac{1}{\sigma_{s}^{2}}+\frac{\alpha_{t|s}^{2}}{\sigma_{t|s}^{2}}\right)^{-1}=\frac{\sigma_{s}^{2}\sigma_{t|s}^{2}}{\sigma_{t}^{2}}.$ (33)
The posterior mean is a convex combination of $\mathbf{G}_{t}$ and $\mathbf{G}_{0}$, with weights determined by their respective noise scales. This provides the optimal denoised estimate of $\mathbf{G}_{s}$ and forms the basis of the reverse sampling process.
A.3 Reverse Denoising Process Approximation
At inference time, the true $\mathbf{G}_{0}$ is unavailable, so we replace it with a model estimate $\hat{\mathbf{G}}_{0}=d_{\theta}(\mathbf{G}_{t},\tilde{\mathbf{G}}_{0},\log\mathrm{SNR}(t))$. This yields the approximate reverse-time transition:
$p(\mathbf{G}_{s}\mid\mathbf{G}_{t})\approx\mathcal{N}(\bar{\mathbf{G}}_{s},\sigma_{Q}^{2}\mathbf{I}),$ (34)
where:
$\bar{\mathbf{G}}_{s}=\frac{\alpha_{t|s}\sigma_{s}^{2}}{\sigma_{t}^{2}}\mathbf{G}_{t}+\frac{\alpha_{s}\sigma_{t|s}^{2}}{\sigma_{t}^{2}}\hat{\mathbf{G}}_{0},$ (35)
$\sigma_{Q}^{2}=\frac{\sigma_{s}^{2}\sigma_{t|s}^{2}}{\sigma_{t}^{2}}.$ (36)
We then sample from this distribution as:
$\mathbf{G}_{s}=\bar{\mathbf{G}}_{s}+\frac{\sigma_{s}\sigma_{t|s}}{\sigma_{t}}\mathbf{G}_{\epsilon},\quad\mathbf{G}_{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (37)
Combining the above expressions, the complete reverse sampling step from $t$ to $s$ becomes:
$\mathbf{G}_{s}=\frac{\alpha_{t|s}\sigma_{s}^{2}}{\sigma_{t}^{2}}\mathbf{G}_{t}+\frac{\alpha_{s}\sigma_{t|s}^{2}}{\sigma_{t}^{2}}\hat{\mathbf{G}}_{0}+\frac{\sigma_{s}\sigma_{t|s}}{\sigma_{t}}\mathbf{G}_{\epsilon}.$ (38)
This closed-form expression enables efficient ancestral sampling in the reverse-time generative process.
A.4 Sampling with Temperature
To control the stochasticity in the sampling process, we introduce a temperature parameter $\tau>0$, which scales the noise component during sampling:
$\mathbf{G}_{s}=\frac{\alpha_{t|s}\sigma_{s}^{2}}{\sigma_{t}^{2}}\mathbf{G}_{t}+\frac{\alpha_{s}\sigma_{t|s}^{2}}{\sigma_{t}^{2}}\hat{\mathbf{G}}_{0}+\tau\cdot\frac{\sigma_{s}\sigma_{t|s}}{\sigma_{t}}\mathbf{G}_{\epsilon}.$ (39)
The temperature parameter $\tau$ modulates the amount of stochasticity in the sampling process. When $\tau=1$, the sampling follows the standard formulation. Lower values of $\tau<1$ reduce the influence of noise, leading to more deterministic and potentially sharper outputs. Conversely, higher values of $\tau>1$ increase the noise contribution, encouraging greater diversity at the cost of higher stochasticity. This flexibility allows practitioners to trade off between sample fidelity and variability based on downstream objectives.
Appendix B Datasets
We conduct experiments on the QM9S dataset [8, 31], which augments the original QM9 [63] molecular dataset with simulated spectra. For each molecule, we extract its structure and corresponding IR, Raman, and UV-Vis spectra, forming a multi-modal spectral condition for our structure elucidation task. The dataset contains 133,885 molecules in total. Following the original setting in DetaNet [8], we adopt the same training, validation, and test split strategy, which partitions the dataset into 90% for training, 5% for validation, and 5% for testing.
Appendix C Metrics
Aiming to generating chemically valid and complete molecules, modeling accurate molecular distributions, and achieving achieve accurate sturcture elucidation from spectra, we elaborately design a series of evaluation metrics to reflect the generation quality.
Specifically, we introduce two categories of evaluation metrics. The first focuses on general molecular generation quality, emphasizing the validity and chemical soundness of the generated structures. The second targets molecular structure elucidation, evaluating the alignment between the predicted molecular structures—derived from spectral data—and the corresponding ground truth structures.
C.1 Basic Metrics for Molecular Generation
In the context of conditional molecule generation, these fundamental evaluation criteria serve as a prerequisite to guarantee that the generated structures are chemically meaningful and sufficiently diverse before evaluating their spectrum-conditioned structural accuracy.
Molecular Structure Validity Evaluation
The primary consideration is the chemical rationality of the generated molecular structures. The Validity metric ensures that the molecular structures conform to basic chemical rules, including chemically plausible valences (e.g., carbon atoms typically form four covalent bonds), and the overall bonding patterns are chemically reasonable (e.g., avoiding unrealistic bonds between atoms, maintaining aromatic delocalization). Building upon this, we introduce the Validity and Complete (V&C) indicator, which not only requires chemical validity but also guarantees that the molecular graph forms a single connected entity instead of fragmented parts. This is particularly important for evaluating the model’s ability to generate fully connected, functional molecules.
To further assess the model’s capacity for innovation, we measure the proportions of molecules satisfying Validity and Unique (V&U) and Validity, Unique, and Novelty (V&U&N) criteria. Uniqueness is determined by canonicalizing molecular representations and removing duplicates, thus ensuring diversity among the generated structures. Novelty is evaluated by comparing the generated molecules against the training set, reflecting the model’s ability to produce genuinely novel chemotypes rather than simply memorizing training data.
Molecular Structure Stability Evaluation
While structural validity is a necessary condition, it may not fully capture the chemical plausibility of a molecule. We additionally introduce a more stringent indicator of molecular stability. Atom Stability assesses whether each atom in the molecule has achieved an appropriate valence coordination, fully considering the influence of formal charges on the permissible bonding patterns. Mol Stability further requires that all atoms in the entire molecule satisfy the stability criteria, thereby providing a more comprehensive assessment of the chemical soundness of the generated molecules.
Distribution-based Structure Evaluation
To evaluate the model’s ability to learn the data distribution of real molecules, we adopt several distribution-based metrics. The Fréchet ChemNet Distance (FCD) measures the similarity between the distributions of generated molecules and reference molecules in a high-dimensional learned feature space. The Similarity to Nearest Neighbor (SNN) metric assesses the representativeness of the generated set by measuring the Tanimoto similarity between each generated molecule and its nearest neighbors in the test set, reflecting both diversity and coverage.
We additionally employ metrics based on molecular structural features: Fragment Similarity (Frag), based on BRICS decomposition, evaluates distributional alignment at the functional group level; while Scaffold Similarity (Scaf), using Bemis–Murcko scaffold analysis, quantifies similarity and diversity at the core-structure level. These complementary metrics jointly reflect the model’s ability to capture both local functional groups and global structural patterns, thereby characterizing the complexity of the generated molecular structures.
3D Geometry-based Structural Evaluation
Beyond chemical rules in 2D representations, assessing the 3D geometric consistency of generated molecules is critical, especially for downstream applications such as molecular docking or molecular structure elucidation from spectra. We therefore introduce a set of geometric evaluation metrics based on fundamental molecular geometry features: bond lengths, bond angles, and dihedral angles.
Bond evaluation compares the distributions of bond lengths in the generated set against those in the reference set (test set molecules), capturing whether typical interatomic distances are preserved. Angle evaluation similarly assesses the distributions of bond angles, reflecting the local spatial arrangements of connected atoms and ring systems. Dihedral evaluation analyzes the distributions of torsional (dihedral) angles, providing insights into conformational consistency and stereochemical plausibility of the generated molecules.
To quantify the similarity between these geometric features of the generated and reference molecules, we employ maximum mean discrepancy (MMD), which measures the distance between two distributions in a kernel space. A lower MMD indicates that the generated molecules more faithfully reproduce realistic 3D geometries.
It is important to note that for all distribution-based and geometric metrics, the test set serves as the reference, while the novelty metric uses the training set as the reference to measure whether generated molecules are unseen. This distinction ensures a clear separation between distributional fidelity and generative innovation.
C.2 Metrics for Molecular Structure Elucidation
In the task of molecular structure elucidation, we introduce a set of precise metrics to evaluate the similarity between the ground-truth target structures and the generated candidate molecules under spectrum-conditioned generation. Here, a molecular structure is represented as a graph $\mathcal{G}=(\mathbf{H},\mathbf{A},\mathbf{X})$, where $\mathbf{H}\in\mathbb{R}^{N\times d_{1}}$ denotes node-level attributes such as atom types and charges, $\mathbf{A}\in\mathbb{R}^{N\times N\times d_{2}}$ encodes pairwise edge features such as bond existence and bond types, and $\mathbf{X}\in\mathbb{R}^{N\times 3}$ corresponds to the 3D coordinates of the atoms.
Top-$K$ Accuracy
The Top-$K$ Accuracy quantifies the model’s ability to recover the exact target structure among its top-$K$ generated candidates. For each condition characterized by a set of spectra, the model is sampled $K$ times to generate $K$ molecular structure candidates. If any of the $K$ candidates exactly matches the ground-truth structure, the prediction is considered correct. Formally, the metric is defined as:
$\operatorname{ACC@}K=\mathbb{E}_{\mathcal{G}}\left[\mathbbm{1}\left(\exists\,i\in\{1,\ldots,K\},\;\hat{\mathcal{G}}_{i}=\mathcal{G}\right)\right],$ (40)
where $\mathbbm{1}(\cdot)$ is the indicator function, $\mathcal{G}$ denotes the ground-truth molecular graph, and $\hat{\mathcal{G}}_{i}$ denotes the $i$-th generated candidate. This metric reflects the probability that the correct structure appears at least once among the top-$K$ model outputs.
Maximum Common Edge Subgraph (MCES)
To quantify graph-structural overlap, we adopt the Maximum Common Edge Subgraph (MCES) metric:
$\operatorname{MCES}=\mathbb{E}_{\mathcal{G}}\left[\max_{\mathcal{H}\subseteq\mathcal{G},\,\mathcal{H}\subseteq\hat{\mathcal{G}}}|E(\mathcal{H})|\right],$ (41)
where $E(\mathcal{H})$ denotes the set of edges in the common subgraph $\mathcal{H}$. This indicator captures the extent of exact subgraph matching between the target and predicted molecular graphs, with higher values reflecting a greater degree of structural recovery.
Fingerprint-based Similarity Metrics
We adopt several fingerprint-based similarity measures. Denote by $\mathbf{a}_{\mathrm{Morgan}}$ and $\hat{\mathbf{a}}_{\mathrm{Morgan}}$ the binary Morgan fingerprint vectors of the target and predicted molecules, respectively, and by $\mathbf{a}_{\mathrm{MACCS}}$ and $\hat{\mathbf{a}}_{\mathrm{MACCS}}$ their MACCS fingerprints. The Tanimoto Similarity over Morgan fingerprints ($\operatorname{TaniSim}_{\mathrm{Morgan}}$) is defined as:
$\operatorname{TaniSim}_{\mathrm{Morgan}}=\mathbb{E}_{\mathcal{G}}\left[\frac{\left|\mathbf{a}_{\mathrm{Morgan}}\land\hat{\mathbf{a}}_{\mathrm{Morgan}}\right|}{\left|\mathbf{a}_{\mathrm{Morgan}}\lor\hat{\mathbf{a}}_{\mathrm{Morgan}}\right|}\right],$ (42)
where $\land$ and $\lor$ denote bitwise AND and OR over the fingerprint vectors. This score quantifies the overlap in local structural patterns. The Cosine Similarity over Morgan fingerprints ($\operatorname{CosSim}_{\mathrm{Morgan}}$) is defined as:
$\operatorname{CosSim}_{\mathrm{Morgan}}=\mathbb{E}_{\mathcal{G}}\left[\frac{\mathbf{a}_{\mathrm{Morgan}}\cdot\hat{\mathbf{a}}_{\mathrm{Morgan}}}{\|\mathbf{a}_{\mathrm{Morgan}}\|\,\|\hat{\mathbf{a}}_{\mathrm{Morgan}}\|}\right].$ (43)
Likewise, the Tanimoto Similarity over MACCS fingerprints ($\operatorname{TaniSim}_{\mathrm{MACCS}}$) is:
$\operatorname{TaniSim}_{\mathrm{MACCS}}=\mathbb{E}_{\mathcal{G}}\left[\frac{\left|\mathbf{a}_{\mathrm{MACCS}}\land\hat{\mathbf{a}}_{\mathrm{MACCS}}\right|}{\left|\mathbf{a}_{\mathrm{MACCS}}\lor\hat{\mathbf{a}}_{\mathrm{MACCS}}\right|}\right].$ (44)
These fingerprints capture interpretable functional patterns based on a set of predefined substructure keys, complementing the Morgan-based representations.
Fragment-based Similarity (Fraggle)
The Fraggle Similarity metric decomposes $\mathcal{G}$ into a set of chemically meaningful fragments =$\mathcal{F}=\{f_{1},f_{2},\dots,f_{n}\}$ by strategies including double acyclic or exocyclic bond cuts with specified fragment size constraints. For each fragment $f_{i}$, Fraggle computes two Tanimoto similarities over RDKit fingerprints between the target molecule $\mathcal{G}$ and the predicted molecule $\hat{\mathcal{G}}$: one based on the standard RDKit fingerprint over the entire molecule, and another based on a masked RDKit fingerprint where atoms outside $f_{i}$ with a Tversky similarity below 0.8 are replaced with wildcards. The fragment-level score is then taken as the maximum of these two values:
$s_{i}=\max\left(\operatorname{TaniSim}_{\mathrm{RDKit}}(\mathbf{a}_{\mathrm{RDKit}},\hat{\mathbf{a}}_{\mathrm{RDKit}}),\operatorname{TaniSim}_{\mathrm{RDKit}}(\mathbf{a}_{\mathrm{RDKit,mask}},\hat{\mathbf{a}}_{\mathrm{RDKit,mask}})\right).$ (45)
The final Fraggle similarity is defined as:
$\operatorname{FraggleSim}=\mathbb{E}_{\mathcal{G}}\left[\max_{f_{i}\in\mathcal{F}}s_{i}\right].$ (46)
This dual approach captures both global and local matching patterns
Functional Group-based Similarity
Lastly, to evaluate functional group consistency, we define a set of chemically significant functional groups (e.g., alkane, alcohol, amine, carboxylic acid, etc.) described by SMARTS patterns. For a pair of molecules $\mathcal{G}$ (ground truth) and $\hat{\mathcal{G}}$ (prediction), we extract their respective functional group sets $FG(\mathcal{G})$ and $FG(\hat{\mathcal{G}})$. The Functional Group Similarity (FGSim) is then computed as:
$\operatorname{FGSim}=\mathbb{E}_{\mathcal{G}}\left[\frac{\left|FG(\mathcal{G})\cap FG(\hat{\mathcal{G}})\right|}{\left|FG(\mathcal{G})\cup FG(\hat{\mathcal{G}})\right|}\right].$ (47)
This interpretable metric is independent of molecular size, computationally efficient via substructure matching, and well-suited for comparative evaluation.